OpenStudy (dmezzullo):

Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8

OpenStudy (dmezzullo):

@KiArUnsYu

OpenStudy (dmezzullo):

@uri

OpenStudy (dmezzullo):

@breja

OpenStudy (dmezzullo):

@yummydum

OpenStudy (anonymous):

\[f(x) = 2(x + 4)\]\[f(x) = 20\]so we can write these both equal to each other since they both equal \(f(x)\):\[20=2(x+4)\]distribute the \(2\):\[20=2x+8\]now just solve for x :)

OpenStudy (dmezzullo):

48

OpenStudy (dmezzullo):

tyvm

OpenStudy (anonymous):

try that again...you did the opposite :) \[20=2x+8\]subtract \(8\):\[12=2x\]divide by \(2\):\[x=6\]

OpenStudy (anonymous):

ya because 2(20+4)= 48

OpenStudy (anonymous):

@dmezzullo your answer is incorrect, \(x=6\)

OpenStudy (dmezzullo):

ahhhh i seee now

OpenStudy (anonymous):

oh oops idk how to do this then

OpenStudy (anonymous):

breja if you put 48 in for x you're going to get a much larger number than 20:\[f(x)=2(48+8)\]\[f(x)=2(56)\]\[f(x)=112\]clearly that's not 20 :)

OpenStudy (anonymous):

k lol :)